# FedKIM: Adaptive Federated Knowledge Injection into Medical Foundation Models

### v2

@@ Abstract
Foundation models have demonstrated remarkable capabilities in handling diverse modalities and tasks, outperforming conventional artificial intelligence (AI) approaches that are highly task-specific and modality-reliant. In the medical domain, however, the development of comprehensive foundation models is constrained by limited access to diverse modalities and stringent privacy regulations. To address these constraints, this study introduces a novel knowledge injection approach, FedKIM, designed to scale the medical foundation model within a federated learning framework. FedKIM leverages lightweight local models to extract healthcare knowledge from private data and integrates this knowledge into a centralized foundation model using a designed adaptive Multitask Multimodal Mixture Of Experts (M3OE) module. This method not only preserves privacy but also enhances the model’s ability to handle complex medical tasks involving multiple modalities. Our extensive experiments across twelve tasks in seven modalities demonstrate the effectiveness of FedKIM in various settings, highlighting its potential to scale medical foundation models without direct access to sensitive data.
@@ 1 Introduction
The first limitation prevents training a medical foundation model from scratch in a centralized manner, while the second one exacerbates the challenge of developing a multimodal, multi-task medical foundation model. To overcome these obstacles, a viable solution is to scale existing medical foundation models and infuse them with medical knowledge. Given that medical data is stored on private clients, the federated learning (FL) paradigm McMahan et al. (2017) offers a promising approach, which is a decentralized and collaborative machine learning method where participants do not need to share data directly. Although several recent studies on federated foundation models Lu et al. (2023); Chen et al. (2024a) have made progress, they primarily focus on enhancing services to local clients using existing foundation models. Importantly, none have specifically tackled the challenge of injecting novel medical knowledge into existing medical foundation models in a federated manner.
@@ 2.1.1 Client Setups
The goal of this work is to scale and enhance the predictive ability of medical large language models (LLMs) by incorporating medical knowledge from private client data in a federated manner. To achieve this, we employ $N$ clients, each representing a hospital or a medical institute holding private medical data $\mathcal{D}_{n}$. We assume that the private dataset $\mathcal{D}_{n}$ contains all medical modalities $\{\mathcal{M}_{1},\cdots,\mathcal{M}_{M}\}$ and can perform all tasks $\{\mathcal{T}_{1},\cdots,\mathcal{T}_{T}\}$. Each client trains a model $f_{n}=[\text{ENC}_{n}();\text{DEC}_{n}()]$ using the data $\mathcal{D}_{n}$, where $\text{ENC}_{n}()$ is the set of multimodal encoders and $\text{DEC}_{n}()$ is the set of multi-task decoders/predictors. Thus, the model parameters $\boldsymbol{\theta}_{n}$ of $f_{n}$ can be divided into $\boldsymbol{\theta}_{n}^{enc}$ for the encoder and $\boldsymbol{\theta}_{n}^{dec}$ for the decoder, which will be further upload to the server.
@@ 2.3.1 Knowledge Aggregation
where $\textit{f}_{FL}$ can be flexibly replaced with any federated learning methods, such as personalized FL methods Jiang et al. (2019); T Dinh et al. (2020), differential privacy-based FL methods Hu et al. (2020); El Ouadrhiri and Abdelhadi (2022), or adaptive FL methods Reddi et al. (2020); Wang et al. (2022b, a).
@@ 4.3 Ablation Study
We conduct an ablation study on the COVID-19 detection task to assess the impact of each module within our proposed FedKIM. Retaining all other modules as in the main experiments, we explore the following variant settings: (1) FedKIMpub: Instead of utilizing knowledge from private datasets $\mathcal{D}_{n}$, this configuration solely leverages public dataset $\mathcal{D}_{p}$ for centralized training. Consequently, the federated training module discussed in Section 2.2 is excluded, with the encoder $\boldsymbol{\theta}_{e}$ updated exclusively through public training as detailed in Section 2.3. (2) FedKIMT: This variant omits the task description module that guides the expert selection process, testing the importance of task-specific information in routing the mixture of experts. (3) FedKIMM: Similarly, we remove the modality description module to examine its influence on expert selection.
Table 4: Ablation study results.
@@
The results of the ablation study are presented in Table 4. They indicate that each component significantly enhances FedKIM’s performance. Specifically, a substantial decline in performance with FedKIMpub highlights the crucial role of knowledge injected from local clients through federated learning. This locally enriched encoder allows the medical foundation model to better adjust to unseen modalities, thereby enhancing its effectiveness compared to models trained without this knowledge. Moreover, the absence of task or modality descriptions diminishes FedKIM’s ability to manage specific tasks through multi-task training, validating the design of the M3OE module. This module equips FedKIM to effectively navigate complex healthcare scenarios that involve diverse tasks and modalities. In summary, the synergistic integration of local knowledge, along with the task and modality description modules, crucially bolsters the performance of our proposed FedKIM.
@@ Appendix B Details of Validation Tasks
Table 6: Single task fine-tuning evaluation.
Task Method LLM FedAvg FedProx
Metric MMedLM-2 FedPlug FedPlugL FedKIM FedPlug FedPlugL FedKIM
COVID-19 Detection Accuracy 92.34 99.56 99.57 84.16 95.66 98.91
Precision 93.59 98.87 99.15 77.27 84.18 100.00
Recall 74.92 99.43 99.15 53.27 98.68 95.73
F1 79.15 99.14 99.15 63.07 90.85 97.82
Lung Opacity Detection Accuracy 89.42 51.69 94.93 91.23 90.90 93.63
Precision 84.69 51.74 93.60 87.76 88.06 92.55
Recall 97.16 99.79 96.85 96.52 95.37 95.37
F1 90.50 68.10 95.19 91.93 91.57 93.94
ECG Abnormal Detection Accuracy 43.15 48.79 58.46 45.25 50.80 58.00
Precision 56.97 65.02 58.46 60.85 58.47 58.34
Recall 11.22 26.82 100.00 17.80 54.67 98.51
F1 18.74 37.98 73.78 27.55 56.51 73.28
Mortality Prediction Accuracy 84.11 91.67 64.16 82.41 90.98 57.38
Precision 16.35 43.24 12.86 13.87 40.68 13.25
Recall 21.43 14.91 56.21 16.64 14.91 72.98
F1 18.55 22.18 20.94 15.13 21.82 22.42

### v3

@@ Abstract
Foundation models have demonstrated remarkable capabilities in handling diverse modalities and tasks, outperforming conventional artificial intelligence (AI) approaches that are highly task-specific and modality-reliant. In the medical domain, however, the development of comprehensive foundation models is constrained by limited access to diverse modalities and stringent privacy regulations. To address these constraints, this study introduces a novel knowledge injection approach, FedKIM, designed to scale the medical foundation model within a federated learning framework. FedKIM leverages lightweight local models to extract healthcare knowledge from private data and integrates this knowledge into a centralized foundation model using a designed adaptive Multitask Multimodal Mixture Of Experts (M3OE) module. This method not only preserves privacy but also enhances the model’s ability to handle complex medical tasks involving multiple modalities. Our extensive experiments across twelve tasks in seven modalities demonstrate the effectiveness of FedKIM in various settings, highlighting its potential to scale medical foundation models without direct access to sensitive data. Source codes are available at  https://github.com/XiaochenWang-PSU/FedKIM.
@@ 1 Introduction
The first limitation prevents training a medical foundation model from scratch in a centralized manner, while the second one exacerbates the challenge of developing a multimodal, multi-task medical foundation model. To overcome these obstacles, a viable solution is to scale existing medical foundation models and infuse them with medical knowledge. Given that medical data is stored on private clients, the federated learning (FL) paradigm McMahan et al. (2017); Che et al. (2023); Zhou et al. (2022) offers a promising approach in the medical domain Wang et al. (2022a), which is a decentralized and collaborative machine learning method where participants do not need to share data directly. Although several recent studies on federated foundation models Lu et al. (2023); Chen et al. (2024a) have made progress, they primarily focus on enhancing services to local clients using existing foundation models. Importantly, none have specifically tackled the challenge of injecting novel medical knowledge into existing medical foundation models in a federated manner.
@@ 2.1.1 Client Setups
The goal of this work is to scale and enhance the predictive ability of medical large language models (LLMs) by incorporating medical knowledge from private client data in a federated manner. To achieve this, we employ $N$ clients, each representing a hospital or a medical institute holding private medical data $\mathcal{D}_{n}$. We assume that the private dataset $\mathcal{D}_{n}$ contains all medical modalities $\{\mathcal{M}_{1},\cdots,\mathcal{M}_{M}\}$ and can perform all tasks $\{\mathcal{T}_{1},\cdots,\mathcal{T}_{T}\}$. Each client trains a model $f_{n}=[\text{ENC}_{n}();\text{DEC}_{n}()]$ using the data $\mathcal{D}_{n}$, where $\text{ENC}_{n}()$ is the set of multimodal encoders and $\text{DEC}_{n}()$ is the set of multi-task decoders/predictors. Thus, the model parameters $\boldsymbol{\theta}_{n}$ of $f_{n}$ can be divided into $\boldsymbol{\theta}_{n}^{enc}$ for the encoder and $\boldsymbol{\theta}_{n}^{dec}$ for the decoder, which will be further uploaded to the server.
@@ 2.3.1 Knowledge Aggregation
where $\textit{f}_{FL}$ can be flexibly replaced with any federated learning methods, such as personalized FL methods Jiang et al. (2019); T Dinh et al. (2020), differential privacy-based FL methods Hu et al. (2020); El Ouadrhiri and Abdelhadi (2022), or adaptive FL methods Reddi et al. (2020); Wang et al. (2022c, b).
@@ 4.3 Ablation Study
We conduct ablation studies on the COVID-19 detection and enlarged cardiomediastinum detection tasks to assess the impact of each module within our proposed FedKIMin both fine-tuning and zero-shot settings. Retaining all other modules as in the main experiments, we explore the following variant settings: (1) FedKIMpub: Instead of utilizing knowledge from private datasets $\mathcal{D}_{n}$, this configuration solely leverages public dataset $\mathcal{D}_{p}$ for centralized training. Consequently, the federated training module discussed in Section 2.2 is excluded, with the encoder $\boldsymbol{\theta}_{e}$ updated exclusively through public training as detailed in Section 2.3. (2) FedKIMT: This variant omits the task description module that guides the expert selection process, testing the importance of task-specific information in routing the mixture of experts. (3) FedKIMM: Similarly, we remove the modality description module to examine its influence on expert selection.
Table 4: Ablation study results in the fine-tuning setting.
@@
Table 5: Ablation study results in the zero-shot setting.
Setting Accuracy Precision Recall F1
FedKIMpub 59.82 54.43 84.40 66.19
FedKIMT 61.11 55.80 79.82 65.66
FedKIMM 60.26 66.00 30.28 41.51
FedKIM 61.54 55.13 93.58 69.39
The results of the ablation studies are presented in Table 4 and Table 5. They indicate that each component significantly enhances FedKIM’s performance. Specifically, a substantial decline in performance with FedKIMpub highlights the crucial role of knowledge injected from local clients through federated learning. This locally enriched encoder allows the medical foundation model to better adjust to unseen modalities, thereby enhancing its effectiveness compared to models trained without this knowledge. Moreover, the absence of task or modality descriptions diminishes FedKIM’s ability to manage specific tasks through multi-task training, validating the design of the M3OE module. This module equips FedKIM to effectively navigate complex healthcare scenarios that involve diverse tasks and modalities. In summary, the synergistic integration of local knowledge, along with the task and modality description modules, crucially bolsters the performance of our proposed FedKIM.
@@ 7 Limitations
Acknowledgements
This work is partially supported by the National Science Foundation under Grant No. 2238275 and 2348541 and the National Institutes of Health under Grant No. R01AG077016.